OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 2x2 – 28x + 84 = 0

OpenStudy (hoblos):

2x2 – 28x + 84 = 0 x2 - 14x +42 =0 x2 -14x + 49 - 49 = -42 (x-7)^2 = 7

OpenStudy (anonymous):

thank you for the help

OpenStudy (anonymous):

\[ 2x^2 – 28x + 84 = 0 \implies x^2-14x+42 = 0 \implies (x-7)^2=7\]

OpenStudy (anonymous):

thank you to

OpenStudy (hoblos):

yw